OpenStudy (anonymous):

-8,-7,0,4,is eometric arithmetic or neither

OpenStudy (anonymous):

for it to be a arithmetic series -7+8=4-0 1=4which is not possible so it is not arithmetic for geometric seris -7/-8=0/4 7/8=0 which is again not possible so answer is neither

OpenStudy (anonymous):

need to know how geometric seqences go from a negative to a positive